OpenStudy (anonymous):

Jane gave the following explanation for factoring the trinomial x2 + 5x + 6: “I need to find factors of 6 that add up to 5. 6 and –1 multiply to 6 and add up to 5 so my factors are (x + 6)(x – 1).” Are Jane’s factors correct? If not, explain where she made her mistake. What could Jane have done to check the accuracy of her factors before submitting her Module 3 test?

OpenStudy (anonymous):

Not correct. 6 and -1 multiply to -6 Expand out

OpenStudy (anonymous):

(x+2)(x+3)

OpenStudy (anonymous):

okay so expand out is for what?

OpenStudy (anonymous):

to see that she was wrong.

OpenStudy (anonymous):

\[ (x-1) (x+6)=x^2+5 x-6 \]

OpenStudy (anonymous):

kindof confused

OpenStudy (anonymous):

confused about what?

OpenStudy (anonymous):

so jane isnt correct and she made her mistake ......

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

You are done. If you do the 3 steps in my first post.

OpenStudy (anonymous):

What could Jane have done to check the accuracy of her factors before submitting her Module 3 test?

OpenStudy (anonymous):

\[ (x-1) (x+6)=x^2+5 x-6 \]

OpenStudy (anonymous):

EXPLAIN where she made her mistake.

OpenStudy (anonymous):

Expand out. I have said it 3 times already.

OpenStudy (anonymous):

Bye.

OpenStudy (anonymous):

goodbye